# Insights into the structure and kinematics of a Milky Way-like galaxy

### v2

@@ 5.1 Radial cross-sections of the galactic disc
The first panel of Figure 8 shows that the gaseous spiral arms exhibit pronounced overdensities, with up to six distinct arms identified as local peaks in the distribution, marked by red dashed vertical lines, with their approximate widths [^6] The peaks were identified by finding local maxima in the running median of the gas surface density, and the widths were estimated using the python peak_width function, which measures the full width at half maximum (FWHM) of each peak. indicated by vertical red shaded areas. These arm positions are in good agreement with the spiral ridges used in Section 4.2 and defined in Appendix C. These spiral arms are overlaid on the subsequent panels, which display the velocity residuals distributions: tangential (blue) and radial (green).
@@
From the tangential velocity residuals (second panel of Figure 8), we can see that the position of the spiral arms seem to be mostly associated areas of positive slope - i.e. where gas is moving slower in the inner side of the arm and faster in the outer side, effectively decreasing the differential rotation (and shear) within the arm. There are exceptions but typically related to lower gas surface density arm ridges. We can see that absolute values of these velocity residuals are larger ($\sim 10-15$ km s-1) closer to the centre, where the spirals are also stronger and better defined, and these deviations decrease with distance.
The third panel of Figure 8 shows the radial velocity component as a function of galactocentric radius. In this frame, a downward slope indicates regions where gas is radially converging, regardless of whether there is a change in sign from outward to inward motion, though the most significant convergence typically corresponds to a sign change. Conversely, an upward (positive) slope suggests divergence in the gas, and a flat slope indicates stable regions, where the gas is neither radially converging nor diverging, regardless of the direction or magnitude of the overall flow. Notably, there is a drop in velocities at the end of the bar region, indicating converging gas. The magnitude of this change ($\sim 40$ km s-1) is much greater than the change in tangential residuals related to the bar ($\sim 15$ km s-1), and it is at the end of this region of strongly converging gas that we find the strongest spiral arms in the gas surface density plot. For larger galactocentric distances along this slice, most gas exhibits relatively stable outwards motions, with the gaseous spiral arms appearing in converging regions where the slope in $\Delta V_{R}$ is negative, and in particular, where it shifts from positive to negative values. The larger radial velocity gradients across the identified spiral arms are of the order of $\sim 12-15$ km s-1 kpc-1. Conversely, clearly diverging regions (with positive slope) are typical of the inter-arm regions, although they have typically less steep slopes than the sharp changes found in spiral arms.
@@
Our analysis suggests that the position and strength of spiral arms strongly correlates with distinct velocity gradients in radial and tangential velocities. We also find that the strongest gaseous spiral arms, characterised by strongly converging gas, are mostly found in the inner galaxy, where the velocity changes are sharper, and partly driven (and sustained) by the bar. These results align with the findings of Urquhart2014, who highlight the prominence of the Sagittarius arm as a major gaseous feature in the Milky Way, based on observations of young massive stars from the Red MSX Source (RMS) survey. It remains to be seen, however, how persistent these velocity patterns are throughout time, which we explore in the following section.
@@ 5.2 Time-evolution of spiral arm segments
We use the gaseous spiral ridge points (see Appendix C) to trace the underlying skeleton of the gaseous spiral structure, and follow the evolution of a selected spiral arm segment (see Fig. 9) over a period of $\sim 20$ Myrs. This particular region was chosen as an example of an arm with a section that is actively growing, and a section that breaks within this time frame (the time evolution of this spiral arm segment [^7] A video of the time-evolution of the velocity fields for the entire galactic disc can be found in the FFOGG project website: https://ffogg.github.io can be seen in Appendix F, Fig. 20).
We investigate the evolution of physical and kinematic conditions of the arm by tracking key quantities along the spiral ridge, namely the gas volume density ($\rho$), radial and tangential velocity residuals ($\Delta V_{\rm R}$ and $\Delta(-V_{\phi})$), and velocity divergence ($\nabla\cdot\mathbf{V}$). In addition, we also look at the fraction of dense gas ($f_{\rm dense}$) as the percentage of mass above a given volume density threshold (which we take to be $700\penalty 10000\ \mathrm{cm}^{-3}$ as the minimum density for gas with star-forming potential, in line with what is used for the insertion of sinks/star particles in similar galaxy-scale simulations, see e.g. dobbs2015; tress2020simulations). The results are presented in Figure 10, where each column shows a different time (spaced by 5 Myrs). The two sections of the spiral arm which are growing (Section 1) and dispersing (Section 2) are shown as shaded grey areas.
@@
Focusing on the arm section that is actively growing (Section 1), we can see that is starts with low densities, and mildly converging motions, and that as time progresses it becomes more convergent and densities increase considerably. This particular section exhibits a lower tangential velocity than the surrounding areas, which means the gas is slowing down in its rotation in this area, and therefore contributing to the gathering of material. This section also exhibits a very mild change of radial velocities along it, but given that the arm segment is near-perpendicular to the radial direction, this suggests this segment has a coherent “bulk motion”, with these radial velocity gradients only mildly reshaping the structure over time. That said, over the period of 20 Myr depicted here, this section evolves from having a bulk motion of mostly outward motions of $\sim 20$ km s-1, to inward motions of $\sim-10$ km s-1, highlighting the rapidly evolving nature of these kinematic patterns. In Fig. 20 we can further see that this particular section of the spiral arm ridge sits at the interface between an outward and inward motion, thus conducive to its growth.
In contrast, this behaviour is reversed in Section 2, where the arm breaks apart. Initially, the gas in this region has a high density peak and shows convergence, but over time it becomes progressively more divergent, eventually leading to the dissolution of the arm. This transition is also marked by a positive gradient in radial velocity residuals, as well as an increase of the tangential component, such that it is rotating faster in the leading part of this section, and hence the two ends of this section effectively feel a different “bulk motion” which contributes to pulling it apart. This process is accompanied by a redistribution of mass, as the volume density progressively decreases. Figure 20, on the bottom row, shows a more interesting perspective on this evolution: on the right-hand end of the arm ridge, the gas was initially converging (in pink) but then a green (diverging) valley appears and progressively splits the ridge into two fronts (one moving radially inwards, and another outwards).
@@
In summary, our combined analysis reveals a strong connection between the large-scale gas dynamics and the localised evolution of spiral arm segments in our galactic model. On global scales, peaks in the gas surface density trace the spiral arms and coincide with marked deviations in both radial and tangential velocity residuals. Gas tends to decelerate in their circular motion inside the arms, while radial velocity patterns highlight alternating converging and diverging flows at spiral arm locations. These large-scale features are consistent with previous N-body simulation results. For example, Kawata2014 found that gas around spiral arms exhibits sharp velocity variations, and baba2015dynamics showed that spiral arms form through swing amplification and dissolve due to dynamical forces. This is also in line with recent Gaia observations by Natsuki2024, which show that the Perseus arm appears to be dissolving as stars escape the arm, while the Outer arm is likely growing with stars converging, both behaviours closely linked to the observed velocity fields. Our results further illustrate the inter-linked and transient nature of spiral structures and gas dynamics. While a spiral pattern is always visible and distinct in the gas, any given spiral segment only survives as a single coherent feature for a few tens of Myrs, and this timescale is dictated by the amplitude of the radial motions present in the gas. Indeed, if this is the dominant mechanism by which arms grow and dissolve, then we might be able to infer the typical lifetime of spiral arms in the Milky Way, by measuring the amplitude of the gas radial motions. Of course the ability for a given gaseous spiral arm to grow and survive large-scale disruption in these dynamically evolving systems, might change if gas self-gravity and stellar feedback were to be included in our model, and we will explore this in future work.
@@ 6 Summary and Conclusion
In this work, we make use of the isothermal simulation presented in DuranCamacho2024 as Model 4, chosen to provide a good representation of the overall observed Milky Way structure, and conduct an in-depth analysis of the structure and kinematical properties of both the stellar and gaseous components across various regions of that model galaxy. In particular, we investigate the influence of large-scale dynamics on defining the large-scale structures, as well as their potential impact on the agglomeration and disruption of gas, which in turn could set the initial conditions for the star formation processes. Our findings can be summarised as follows:
  • The stellar distribution of our model’s inner galaxy (Section 3.1) can be well described by a boxy/peanut bar structure with a half-length of $\sim$2 kpc and a long bar component with a half-length of 3.1 kpc, aligning well with both analytical models and observational data of the Milky Way (bissantz2002spiral; wegg2015structure; sormani2022stellar; ridley2017nuclear). Nonetheless, the model exhibits a thinner surface density distribution within the inner 6 kpc, with the largest discrepancies in the vertical direction.
  • The kinematic pattern of our model’s inner galaxy (Section 3.2) shows a radial velocity quadrupole pattern, and a tangential velocity deceleration region along the bar, both with similar extent and amplitude to the observations from Gaia (e.g. GaiaDR32021; drimmel2023).
@@
  • We analysed the stellar kinematics around the solar neighbourhood (Section 4.2) and compare to the observed stellar velocity distribution from Gaia DR3 (e.g. khanna2023measuring; drimmel2023). We find our model to be largely in agreement with the observations (albeit with slightly larger amplitudes in the velocity residuals), but we find that there is not always a unique correlation between specific changes in the velocity of the stars and the position of the spiral arms (may those be the gas or stellar arms). However, we find that the gas experiences much shaper kinematic changes, in both radial and tangential directions, which correlate better with the position of the gaseous spiral arms.
@@
The maps for the top-down surface density, velocity and lv projection of our best model have been made publicly available through the publication (DuranCamacho2024), and can be found at the Following the Flow Of Gas in Galaxies (FFOGG) project website (https://ffogg.github.io/). Any data not available on the website can be provided upon requests to the authors. With this article, we further make available a video of the time evolution of the spiral structure of our simulation, with specific focus on the evolution of the different kinematical patterns studied here.
@@ Appendix A Inner Galactic Stellar Profiles
In Section 3, we investigate the stellar surface density distribution in our model, and compare it to the work from sormani2022stellar and ridley2017nuclear. Here we provide more details on that comparison, as well as the fitting of the S22-type of profile to our model.
sormani2022stellar presents an analytical model tailored to describe the stellar profile of the inner Galaxy, based on the N-body model from portail2017dynamical, which had been developed to match observational data (e.g. wegg2013mapping; wegg2015structure). Their global density profile (Eq. 1) includes 4 components: a boxy/peanut region in the centre which is described by the components bar1,2, a long bar described by bar3, and an axisymmetric disc.
@@
Figure 11: Stellar surface density along the $x$ (left), $y$ (middle), and $z$ (right) axes of our model compared to the sormani2022stellar analytic model (S22). Top row shows the comparison between our model (red-dotted line) and the S22 model (yellow-crossed line). The different coloured-dashed lines represent the various galactic components of the S22 model: disc in dark blue, bar1{1} in pink, bar2{2} in green, and bar3 or long bar in light blue. Bottom row includes the overall best fit of our model (black solid line) along with its components, represented by the same colour-coded dashed lines as in figures in the top row. Left-hand and middle panels correspond to the $x$ and $y$ axes in the ($x$,$y$) top-view map, while right-hand panels corresponds to the $z$-axis in the ($x$,$z$) map.
@@
Table 1: Best-fitting parameters for our numerical model, compared to the analytical model by sormani2022stellar. S22 values are shown only when they differ from this work.
@@
We also compare our model with the hydrodynamical simulation presented in ridley2017nuclear, which simulate the dynamics of CMZ, and is based on observations of NH3 (1,1) emission from the HOPS survey and an improved potential from McMillan2017. The R17 model includes 3-dimensional profiles for three components: bar, bulge, disc and halo. For the purpose of comparing the R17 model to our model’s stellar profile in the galactic centre alone, we exclude the halo component of R17, and confine our analysis to the inner 3 kpc, where that component is negligible.
@@
Figure 14: Cylindrical enclosed mass as a function of galactocentric radius for the analytical model from (sormani2022stellar) (cross-dashed lines), and our model (dot-solid lines). The different colours refer to the different $z$-slices: purple for the mid-slice ($0<|z|<0.25$ kpc), blue for the upper-slice ($0.25<|z|<0.5$ kpc) and green for the top-slice ($0.5<|z|<0.75$ kpc). The position of the two ILRs in our model are shown as green vertical dashed lines.

### v3

@@ 5.1 Radial cross-sections of the galactic disc
The first panel of Figure 8 shows that the gaseous spiral arms exhibit pronounced overdensities, with up to six distinct arms identified as local peaks in the distribution at $R\simeq 4.0,4.2,5.8,7.0,8.5,$ and $11.1$ kpc, marked by red dashed vertical lines, with their approximate widths [^6] The peaks were identified by finding local maxima in the running median of the gas surface density, and the widths were estimated using the python peak_width function, which measures the full width at half maximum (FWHM) of each peak. indicated by vertical red shaded areas. These peak radii are in good agreement with the positions where the extracted gaseous spiral ridges intersect this line-of-sight (Appendix C) and are consistent with the ridges used in Sect. 4.2. For reference, the same arm intervals are overplotted in the lower panels, which show the tangential (middle) and radial (bottom) velocity residuals along the same slice.
@@
In the tangential residual profile (Fig. 8, middle panel), the position of the spiral arms are mostly associated with a positive slope, i.e. $\Delta(-V_{\phi})$ increases across the arm from its inner to outer edge. This corresponds to gas rotating slower on the inner side of the arm and more rapidly on the outer side, reducing the local differential rotation (shear) within the arm. Deviations from this trend occur mainly for the weakest arm peaks (e.g. at $R\simeq 8.5$). The amplitudes are largest in the inner disc ($|\Delta(-V_{\phi})|\sim 10$–15 km s-1 at $R\lesssim$ 8 kpc), where the spiral features are strongest, and decline at larger radii.
The third panel of Figure 8 shows the radial velocity component as a function of galactocentric radius. In this frame, a downward slope indicates regions where gas is radially converging, regardless of whether there is a change in sign from outward to inward motion, though the most significant convergence typically corresponds to a sign change. Conversely, an upward (positive) slope suggests divergence in the gas, and a flat slope indicates stable regions, where the gas is neither radially converging nor diverging, regardless of the direction or magnitude of the overall flow. A prominent example is the sharp drop in $\Delta V_{R}$ at the end of the bar region (grey shaded area), with an amplitude of $\sim 40$ km s-1. This is substantially larger than the bar-related change in $\Delta(-V_{\phi})$ along the same slice ($\sim 15$ km s-1), and it is at the end of this region of strongly converging gas that we find the strongest spiral arms in the gas surface density plot. At larger radii, most gas exhibits relatively stable outwards motions, with the gaseous spiral arms located where $\Delta V_{R}$ shows locally negative slopes (often where $\Delta V_{R}$ transitions from positive to negative), with typical gradients across the arms of $\sim 12$–15 km s-1 kpc-1. In contrast, inter-arm regions more often show positive slopes (divergence), generally with weaker gradients than those associated with the arm peaks.
@@
Our analysis suggests that the position and strength of spiral arms strongly correlates with distinct velocity gradients in radial and tangential velocities. We also find that the strongest gaseous spiral arms, characterised by strongly converging gas, are mostly found in the inner galaxy, where the velocity changes are sharper. These velocity gradients occur predominantly in the inner disc, within/around the bar corotation radius ($R_{\rm CR}\simeq 6.1$ kpc for this model; see the resonance analysis in Durán-Camacho et al. (2024)), where the bar-driven non-axisymmetric streaming field is expected to be strongest. These results align with the findings of Urquhart et al. (2014), who highlight the prominence of the Sagittarius arm as a major gaseous feature in the Milky Way, based on observations of young massive stars from the Red MSX Source (RMS) survey. It remains to be seen, however, how persistent these velocity patterns are throughout time, which we explore in the following section.
@@ 5.2 Time-evolution of spiral arm segments
We use the gaseous spiral ridge points (see Appendix C) to trace the underlying skeleton of the gaseous spiral structure, and follow the evolution of a selected spiral arm segment (see inset box in Fig. 9) over a period of $\sim 20$ Myrs. This particular region was chosen as an example of an arm with a section that is actively growing, and breaks within this time frame (the time evolution of this spiral arm segment [^7] A video of the time-evolution of the velocity fields for the entire galactic disc can be found in the FFOGG project website: https://ffogg.github.io can be seen in Fig. 20).
We investigate the evolution of physical and kinematic conditions of the arm by tracking key quantities as a function of distance along the spiral ridge, $s$, namely the gas volume density ($\rho$), radial and tangential velocity residuals ($\Delta V_{\rm R}$ and $\Delta(-V_{\phi})$), and velocity divergence ($\nabla\cdot\mathbf{V}$). In addition, we also track the fraction of dense gas ($f_{\rm dense}$), defined as the percentage of mass above a volume density threshold of $700\penalty 10000\ \mathrm{cm}^{-3}$ (in line with sink/star-particle insertion thresholds used in similar galaxy-scale simulations; see e.g. Dobbs, 2015; Treß et al., 2021).
The results are presented in Figure 10, where each column corresponds to a different snapshot (separated by 5 Myr), while the four rows colour-code the ridge points by $\Delta V_{\rm R}$, $\Delta(-V_{\phi})$, $\nabla\cdot\mathbf{V}$, and $f_{\rm dense}$, respectively. Throughout this section we refer to two fixed ridge intervals (shaded in grey in Fig. 10): a growing segment (Section 1) at $s\simeq$ 1.21.21.81.8 kpc and a dispersing segment (Section 2) starting at $s\sim$ $3.1$ kpc.
@@
Focusing first on the growing segment, Fig. 10 shows that Section 1 initially appears comparatively low-density and only weakly convergent, as indicated by small negative values of $\nabla\cdot\mathbf{V}$ (third row), and that as time progresses it becomes more convergent and densities increase considerably.
The kinematic evolution is consistent with this build-up, as this particular section exhibits lower $\Delta(-V_{\phi})$ values (second row) than the surrounding areas, indicating that gas is slowing down in its rotation, and therefore contributing to the gathering of material.
This section also exhibits a very mild change of radial velocities (first row) along it, but given that the arm segment is nearly perpendicular to the radial direction, it suggests that this segment has a coherent “bulk motion”, with these radial velocity gradients only mildly reshaping the structure over time. Over the 20 Myr shown, the typical $\Delta V_{\rm R}$ values within Section 1 shift from predominantly positive (outward) at early times to predominantly negative (inward) at late times, illustrating that the local flow field evolves rapidly even as the ridge grows in density. This behaviour is consistent with Fig. 20, where Section 1 ($x\simeq[-3.3,-0.5]$ kpc and $y\simeq[5.4,5.8]$ kpc) lies close to an interface between outward and inward motions, i.e. a convergence zone in the 2D field. Finally, $f_{\rm dense}$ (fourth row) increases within the same $s$-interval as $\rho$ rises, indicating that growth of the ridge is accompanied by an increasing fraction of gas above the star-forming density threshold.
In contrast, the dispersing segment shows the opposite sequence. Initially, the gas in this region has a high density peak and shows convergence, but over time it becomes progressively more divergent, eventually leading to the dissolution of the arm. This transition is also marked by a positive gradient in radial velocity residuals, as well as an increase of the tangential component, such that it is rotating faster in the leading part of this section, and hence the two ends of this section effectively feel a different “bulk motion” which contributes to pulling it apart. This process is accompanied by a redistribution of mass, as the volume density progressively decreases. Consistently, $f_{\rm dense}$ decreases as the ridge disperses. The 2D perspective shown in Fig. 20 supports this interpretation: at the location corresponding to Section 2 ($x\simeq[-1.6,0.1]$ kpc), the converging region present at early times is progressively replaced by a diverging “valley”, which splits the ridge into two fronts that subsequently move radially in opposite directions.
@@
In summary, our combined analysis reveals a strong connection between the large-scale gas dynamics and the localised evolution of spiral arm segments in our galactic model. On global scales, peaks in the gas surface density trace the spiral arms and coincide with marked deviations in both radial and tangential velocity residuals. Gas tends to decelerate in their circular motion inside the arms, while radial velocity patterns highlight alternating converging and diverging flows at spiral arm locations. These large-scale features are consistent with previous N-body simulation results. For example, Kawata et al. (2014) found that gas around spiral arms exhibits sharp velocity variations, and Baba et al. (2013) showed that spiral arms form through swing amplification and dissolve due to dynamical forces. This is also in line with recent Gaia observations by Funakoshi et al. (2024), which show that the Perseus arm appears to be dissolving as stars escape the arm, while the Outer arm is likely growing with stars converging, both behaviours closely linked to the observed velocity fields. Our results further illustrate the inter-linked and transient nature of spiral structures and gas dynamics. While a spiral pattern is always visible and distinct in the gas, any given spiral segment only survives as a single coherent feature for a few tens of Myrs, and this timescale is dictated by the amplitude of the radial motions present in the gas. Indeed, if this is the dominant mechanism by which arms grow and dissolve, then we might be able to infer the typical lifetime of spiral arms in the Milky Way, by measuring the amplitude of the gas radial motions. Of course the ability for a given gaseous spiral arm to grow and survive large-scale disruption in these dynamically evolving systems, might change if gas self-gravity and stellar feedback were to be included in our model, and we will explore this in future work.
@@ 6 Summary and conclusion
In this work, we make use of the isothermal simulation presented in Durán-Camacho et al. (2024) as Model 4, chosen to provide a good representation of the overall observed Milky Way structure, and conduct an in-depth analysis of the structure and kinematical properties of both the stellar and gaseous components across various regions of that model galaxy. In particular, we investigate the influence of large-scale dynamics on defining the large-scale structures, as well as their potential impact on the agglomeration and disruption of gas, which in turn could set the initial conditions for the star formation processes. Our findings can be summarised as follows:
  • The stellar distribution of our model’s inner galaxy (Section 3.1) can be well described by a boxy/peanut bar structure with a half-length of $\sim$2 kpc and a long bar component with a half-length of 3.1 kpc, aligning well with both analytical models and observational data of the Milky Way (Bissantz and Gerhard, 2002; Wegg et al., 2015; Sormani et al., 2022; Ridley et al., 2017). Nonetheless, the model exhibits a thinner surface density distribution within the inner 6 kpc, with the largest discrepancies in the vertical direction.
  • The kinematic pattern of our model’s inner galaxy (Section 3.2) shows a radial velocity quadrupole pattern, and a tangential velocity deceleration region along the bar, both with similar extent and amplitude to the observations from Gaia (e.g. Gaia Collaboration et al., 2021b, 2023).
@@
  • We analysed the stellar kinematics around the solar neighbourhood (Section 4.2) and compare to the observed stellar velocity distribution from Gaia DR3 (e.g. Khanna et al., 2023; Gaia Collaboration et al., 2023). We find our model to be largely in agreement with the observations (albeit with slightly larger amplitudes in the velocity residuals), but we find that there is not always a unique correlation between specific changes in the velocity of the stars and the position of the spiral arms (may those be the gas or stellar arms). However, we find that the gas experiences much shaper kinematic changes, in both radial and tangential directions, which correlate better with the position of the gaseous spiral arms.
@@
The maps for the top-down surface density, velocity and lv projection of our best model have been made publicly available through the publication (Durán-Camacho et al., 2024), and can be found at the Following the Flow Of Gas in Galaxies (FFOGG) project website (https://ffogg.github.io/). Any data not available on the website can be provided upon requests to the authors. With this article, we further make available a video of the time evolution of the spiral structure of our simulation, with specific focus on the evolution of the different kinematical patterns studied here.
@@ Appendix A Inner Galactic stellar profiles
In Section 3, we investigate the stellar surface density distribution in our model, and compare it to the work from Sormani et al. (2022, S22) and Ridley et al. (2017, R17). Here we provide more details on that comparison, as well as the fitting of the S22-type of profile to our model.
Sormani et al. (2022) presents an analytical model tailored to describe the stellar profile of the inner Galaxy, based on the N-body model from Portail et al. (2017), which had been developed to match observational data (e.g. Wegg and Gerhard, 2013; Wegg et al., 2015). Their global density profile (Eq. 1) includes 4 components: a boxy/peanut region in the centre which is described by the components bar1,2, a long bar described by bar3, and an axisymmetric disc.
@@
Figure 11: Stellar surface density along the $x$ (left), $y$ (middle), and $z$ (right) axes of our model compared to the Sormani et al. (2022) analytic model (S22). Top row shows the comparison between our model (red-dotted line) and the S22 model (yellow-crossed line). The different coloured-dashed lines represent the various galactic components of the S22 model: disc in dark blue, bar1{1} in pink, bar2{2} in green, and bar3 or long bar in light blue. Bottom row includes the overall best fit of our model (black solid line) along with its components, represented by the same colour-coded dashed lines as in figures in the top row. Left-hand and middle panels correspond to the $x$ and $y$ axes in the ($x$,$y$) top-view map, while right-hand panels corresponds to the $z$-axis in the ($x$,$z$) map.
@@
Table 1: Best-fitting parameters for our numerical model. [^8] Values are compared to the analytical model by Sormani et al. (2022). S22 values are shown only when they differ from this work.
@@
We also compare our model with the hydrodynamical simulation presented in Ridley et al. (2017), which simulate the dynamics of CMZ, and is based on observations of NH3 (1,1) emission from the HOPS survey and an improved potential from McMillan (2017). The R17 model includes 3-dimensional profiles for three components: bar, bulge, disc and halo. For the purpose of comparing the R17 model to our model’s stellar profile in the galactic centre alone, we exclude the halo component of R17, and confine our analysis to the inner 3 kpc, where that component is negligible.
@@
Figure 14: Cylindrical enclosed mass as a function of galactocentric radius for the analytical model from (Sormani et al., 2022) (cross-dashed lines), and our model (dot-solid lines). The different colours refer to the different $z$-slices: purple for the mid-slice ($0<|z|<0.25$ kpc), blue for the upper-slice ($0.25<|z|<0.5$ kpc) and green for the top-slice ($0.5<|z|<0.75$ kpc). The position of the two ILRs in our model are shown as green vertical dashed lines.